Sarah Meeham blends coffee for​ Tasti-Delight. She needs to prepare 150 pounds of blended coffee beans selling for ​$4.47 per pound. She plans to do this by blending together a​ high-quality bean costing ​$5.00 per pound and a cheaper bean at ​$3.00 per pound. To the nearest​ pound, find how much​ high-quality coffee bean and how much cheaper coffee bean she should blend.
Found 2 solutions by stanbon, addingup:
Answer by stanbon(75887)

Sarah Meeham blends coffee for​ Tasti-Delight. She needs to prepare 150 pounds of blended coffee beans selling for ​$4.47 per pound. She plans to do this by blending together a​ high-quality bean costing ​$5.00 per pound and a cheaper bean at ​$3.00 per pound. To the nearest​ pound, find how much​ high-quality coffee bean and how much cheaper coffee bean she should blend.
-----
Equations:
Quantity:: H + L = 150 lbs
Value::: 5H + 3L = 4.47*150
------------------------------
Modify to solve for "H"::
3H + 3L = 3*150
5H + 3L = 4.47*150
------------------------------
Subtract and solve for "H"
2H = 1.47*150
2H = 110.25
H = 55.125 lbs (amt. of high-quality beans needed)
Solve for "L"::
L = 150-55.125 = 94.875 lb (amt. of lower quality beans needed)
----------
Cheers,
Stan H.
------------

Answer by addingup(3677)

150(4.47) <-- this is the desired outcome, 150 lbs of 4.47 blend
By mixing x lbs of 5.00 with 150-x lbs of 3.00
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
5.00x + 3.00(150-x) = 150(4.47)
5.00x + 450 - 3.00x = 670.50
2.00x = 220.50
x = 110.25
Sarah needs 110.25 lbs of the $5.00 beans and 150-110.25 = 39.75 lbs of the $3.00 beans